OpenStudy (anonymous):

Solve: 4(x − 1) = 2(x + 2) + 6

OpenStudy (shaik0124):

4x-4=2x+4+6 4x-2x=10+4 2x=14 x=14/2 x=7

OpenStudy (shaik0124):

@owenmisa123

OpenStudy (anonymous):

thanx